If a spring of stiffness ‘k’ is cut into parts ‘A’ and ‘B’ of length ℓA : ℓB = 2 : 3, then the stiffness of spring ‘A’ is given by
(a) 3k/5
(b) 2k/5
(c) k
(d) 5/2k

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
(d) Here, ℓA = 2ℓ/5, ℓB = (3ℓ/5)
kℓ = kAA = kBB kℓ = kA(2ℓ/5
kA = 5k/2 \Rightarrow kB =5k/3